OpenStudy (anonymous):

i have a couple of questions Choose the system of equations which matches the following graph.

OpenStudy (anonymous):

alright let's see

OpenStudy (anonymous):

ill upload it but how do i upload it

OpenStudy (anonymous):

"attach file" below

OpenStudy (anonymous):

do you have word

OpenStudy (anonymous):

or adobe

OpenStudy (anonymous):

well first of all, u move y to the left side so as u have y=....

OpenStudy (anonymous):

can you just tell me the answers because these problems i been having trouble on

OpenStudy (anonymous):

well look,the letters beside x like A and D can be solved for like this,u take 2 points from the graph,1(x1,y1) and 2(x2,y2) and then,A= y1-y2/x1-x2 then u put that in ur equation,isolate y,replace it by a value of y in ur graph and its value of x in ur x and solve

OpenStudy (anonymous):

can you just tell me the answers plz cause i have to finish in 10mins

OpenStudy (anonymous):

wow 10min! i can't sorry,it's already 5 am here,sorry

OpenStudy (anonymous):

why cant you help me

OpenStudy (anonymous):

well in the last part, -89 can't be put as a value of x cuz u can't have a negative number under the root

OpenStudy (anonymous):

for x= 51, 3(7)+7=28 21+7=28 correct

OpenStudy (anonymous):

for 9 u get 3.sqr of 7+7=28 ,u can't add up these two on the left side so...

OpenStudy (anonymous):

same story for the rest,apart 51

OpenStudy (anonymous):

hope i helped a bit,it's the end of summer school u know it's been a while i didn't practice math

OpenStudy (anonymous):

so whats the answer 51 for last one